I'm 46 years old and am trying to relearn algebra on my own. Any help is greatly appreciated.
How do I solve the following:
In baseball a player's batting average is the ratio of hits to the number of times he or she has batted. Thus if the player has an average of .300, it means he or she got 300 hits in 1000 at bats. In order to have an average of .300, how many hits would the batter have to get in 400 times at bat?

Answer by lwsshak3(11628)

In baseball a player's batting average is the ratio of hits to the number of times he or she has batted. Thus if the player has an average of .300, it means he or she got 300 hits in 1000 at bats. In order to have an average of .300, how many hits would the batter have to get in 400 times at bat?
..
This is a ratio/proportion problem
let x=hits required to get a batting average of .300 with 400 times at bat
.300=300/1000=x/400
.3=x/400
cross-multiply
x=.3(400)=120 times at bat
check:
120/400=.300